ELECTRON SPIN RESONANCE INVESTIGATIONS OF MATRIX ISOLATED METAL CLUSTERS AT 4 K

\underline{BRADFORD M. MESS}, JOHN J. BANISAUKAS, and LON B. KNIGHT, JR., Department of Chemistry, Furman University, Greenville, SC 29613.

High-spin magnesium and calcium neutral clusters and zinc cluster radical cations were isolated in neon matrices at 4 K and studied by ESR spectroscopy. Solid magnesium, calcium, and zinc were vaporized in a resistively heated tantalum Knudsen cell and co-deposited with neon or argon on a copper flat at 4 K. High-spin neutral metal clusters were observed for both magnesium (^{3}Σ) and calcium (^{5}Σ). Experimental values for the fine structure interactions (D-tensor) were determined by fitting the observed fine structure lines to those obtained from an exact diagonalization of the spin Hamiltonian. Due to the large proportion of p-orbital contribution, no hyperfine interactions were resolved. In these experiments, zinc cluster cation radicals were also generated by X-irradiation of matrix isolated zinc clusters. Experimental values of the nuclear hyperfine interactions (A-tensor) for ^{67}Zn_{n}^{+} (n = 1-4) radicals were determined. Small shifts observed in g-values demonstrate some d-orbital contribution to cluster bonding.